OpenStudy (momopeace):

Write 12h^2-31h+20 as the product of two factors. What are the zeros of the related function?

ganeshie8 (ganeshie8):

Split -31 as -16 - 15

OpenStudy (momopeace):

12h^2(-16-15)h+20? How do you rewrite that

OpenStudy (mathstudent55):

Factoring a trinomial of the form \(ax^2 + bx + c\), where \(a\ne 1\) 1. First try to factor a common term out of all factors. 2. Multiply ac together. 3. Find two numbers that multiply to ac and add to b. Call these numbers p and q. 4. Split up the middle term, bx, into px + qx. 5. Factor by grouping.

OpenStudy (mathstudent55):

Step 1. 12, -31, and 20 have no common factors, so we cannot factor out a common factor. Step 2. ac = 12 * 20 = 240 Step 3. We need 2 numbers that multiply to 240 and add to -31. @ganeshie8 alreay supplied the numbers. They are -15 and -16. Step 4. Split up the bh term term into two terms using the numbers -15 and -16. \(12h^2 - 15 h - 16h + 20\) Step 5. Factor by grouping. That means, take out a common factor out of the first two terms. Take out a common factor out of the last two terms. Then take out a common factor (a binomial). \(12h^2 - 15h - 16h + 20\) \(= 3h(4h - 5) - 4(4h - 5) \) \(= (4h - 5)(3h - 4) \)

OpenStudy (momopeace):

Thank yyouu :)

OpenStudy (momopeace):

For the second part of the questions, "what are the zeros of the related function?" What exactly is that asking for?

OpenStudy (momopeace):

@mathstudent55

OpenStudy (mathstudent55):

They want you to set that polynomial equal to zero and solve it as an equation.

OpenStudy (mathstudent55):

Since we already factored it, set each factor equal to zero, and solve it for h. 4h - 5 = 0 or 3h - 4 = 0 Solve both equations above and give the answers separated by the word "or".